The Mead Paper Company produces rolls of paper for cash registers and office equipment. The rolls are cut in widths of 2, 3 and 4 inches. Before cutting, the rolls arrive 10 inches wide. The company must decide how to produce the three different widths efficiently without wasting any paper. The following table shows five cutting alternatives to consider. Number Rolls of Cut alternative 2 in 3 in 4in 0 0 3 0 2 2 0 0 4 0 The minimum requirements for the three products are 1200 units of 2 inch rolls, 1800 units of 3 inch rolls, and 1000 units of 4 inch rolls. If the company wants to minimize the number of 10 inch rolls that must be cut, how many 10 inch rolls will need to be processed on each cutting alternatives? Create the entire linear programming model below and plug into LINDO and interpret AND show the results. (Hint: Let Xi- number of 10 inch rolls of paper processed by cutting alternative i, where i-1,2,3,4,5.)
Answer #1

Solution:

Linear Programming model is as follows

Let Xi = number of 10 inch rolls of paper processed by cutting alternative i, where i ϵ {1,2,3,4,5}

Min X1+X2+X3+X4+X5

s.t.

5X1+0X2+2X3+1X4+3X5 >= 1200

0X1+3X2+2X3+1X4+0X5 >= 1800

0X1+0X2+0X3+1X4+1X5 >= 1000

Xi >= 0

Solution using LINDO is as follows:

Lingo 17.0 Solution Report Lingol Eile Edit Solver Window Help D 레日 @j L Brlo elal.RI庮 ?h@ Solution Report -Lingo1 Lindo Model - Lingo Min Xl+X2+X3+X4+X5 Global optimal solution found Objective value: Infeasibilities: Total solver iterations: Elapsed runtime seconds: 1300.000 0.000000 0.21 LF 5X1+0X2 +2X3+2X4 +3X5 >= 1200 02.1+3X2+2X3+2X4+0X5 >= 1800 Model Class: 021 +0X2 +0X3+2X4+2X5 1000 Total variables: Nonlinear variables: Integer variables: >= Total constraints: Nonlinear constraints: 14 Total nonzeros: Nonlinear nonzeroS: Variable X1 X2 X3 X4 X5 Value 0.000000 200.0000 100.0000 1000.000 0.000000 Reduced Cost 0.1666667 0.000000 0.000000 0.000000 0.000000 Row Slack or Suxplus 1300.000 0.000000 0.000000 0.000000 Dual Price -1.000000 -0.1666667 -0.3333333 -0.5000000

Solution:

X1 = 0

X2 = 200

X3 = 100

X4 = 1000

X5 = 0

I hope this helps if you find any problem. Please comment below. Don't forget to give a thumbs up if you liked it. :)